A heater with a resistance of 100  is connected to a current source with an EMF of 48 V and 

an internal resistance of 1.5 . Find the current power on the heater.


1
Expert's answer
2020-11-19T06:00:55-0500

Current power on the heater is


"P=I^2R,"


where I - current through the heater;

R - resistance of heater.


Current through the heater is


"I=\frac{E}{R+R_{int}},"


where E - EMF of the current source;

Rint - its internal resistance.


So,


"P=\frac{E^2R}{(R+R_{int})^2}=\frac{48^2 \cdot 100}{(100+1.5)^2}=22.36 \space W"


Answer: 22.36 W